(2) (a) What is product differentiation? In what market structure is this an essential feature? Briefly explain, at least, three different ways firms in this industry differentiate their products from those of other firms.

                                                                                                                                                                                                        

(b)Suppose a firm in this industry faces the following demand and total cost functions: 

                                                                                             P = 5900 – 10Q  

                                                                                                          3     2

                                                                                            TC = 2Q - 4Q +140Q + 845.

Determine the profit maximizing output the firm should produce, the market clearing price and the total profit (if any) to be earned.


 

 


1
Expert's answer
2021-12-15T16:20:20-0500

a.) Product differentiation refers to the process of distinguishing a product or service from others, to make it more attractive to a particular target market. The three main categories of product differentiation include the following;




  • Vertical differentiation - Here products are differentiated using price.
  • Horizontal differentiation - Here products differentiation is not associated with the price nor the quality of the product.
  • Mixed differentiation - Here product differentiation is based on a number of several factors.


b). Profit maximizing output is where MR = MC


Derive TR:


TR = P "\times" Q = (5900-10Q)"\times" Q = 5900Q-10Q2

MR = 5900 - 20Q

MC = 2 - 4 + 140


Set MR=MC


5900-20Q=2-4+140


Q=288


Profit maximizing quantity = 288

Price = 5900-10"\times"(288) = 5900-2880 = 3,020




Profit = TR - TC


TR = 3020 "\times" 288 = 869,760

TC = 2 "\times" (288)- 4"\times" (288)+140 "\times" (288)+845


TC = 39,828


Profit = 869,760- 39,828 = 829,932